OpenStudy (spicytaco1):

Molly works at an electronics store. Her last customer purchased a DVD player for $64.76 and five DVDs for $60.19, including taxes. If the customer paid her with two $100 bills, how much change should she give him? A. $135.24 B. $75.05 C. $66.25 D. $139.81

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{2\times 100-(64.76+60.19)\\~\\ =200-124.95=? }\end{align}\)